OpenStudy (anonymous):

In what case is 2+2=5?

OpenStudy (anonymous):

when you read the two wrong.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

via occams razor

OpenStudy (anonymous):

this actually has an answer i just got bored.

OpenStudy (anonymous):

i noe.. thats exactly how i feel.. ..

OpenStudy (anonymous):

My calculus teacher in my junior year of high school proved this but i forgot how

OpenStudy (anonymous):

well 2.2 rounds down to 2 but 2.3+2.3=4.6, which rounds up to 5.

OpenStudy (anonymous):

thats a pretty lame proof.

OpenStudy (anonymous):

oh hush. i dont feel like writing a proof.

OpenStudy (anonymous):

i dont blame yea!

OpenStudy (anonymous):

This is going to bug me now lol. Something with variables and tweaking the problem so it makes it equal 5 but its a bullpellet way to get to it

OpenStudy (anonymous):

that sentence ^^ requires a preety gud ans.

OpenStudy (anonymous):

let a = 0.5, b = 0.5: a^2 = b^2 a^2 - ab = b^2 - ab a(a - b) = -b(a - b) divide both sides by a - b: a = -b thus 0.5 = -0.5 add 0.5 to both sides: thus 1 = 0 add 4 to both sides, thus 5 = 4 thus 5 = 2 + 2

OpenStudy (anonymous):

Thank you jamesm! That's what it was